OpenStudy (anonymous):

Lagrange: f(x,y) = x^2*y^2 9x^2+4y^2 = 36 Find local max and min

OpenStudy (anonymous):

could you explain the answer?

OpenStudy (anonymous):

g(x) = 9x^2 + 4y^2 - 36 so del(f) = λdel(g) delf = 2x*y^2i + 2y*x^2j delg = 18xi + 8yj 2xy^2 = λ18x 2yx^2 = λ8y y^2 =9 λ x^2 = 4λ 9*4λ + 4*9λ - 36 = 0 72λ = 36 λ = 1/2 y^2 = 4.5 x^2 = 2 y = +/- sqrt(4.5) x = +/- sqrt(2)

OpenStudy (anonymous):

CoolSector Saves the Day!

OpenStudy (anonymous):

hope im right lol

OpenStudy (anonymous):

hee yeahh o.o but what about the local min and maX?

OpenStudy (anonymous):

oh you have to plug the values and see which is bigger

OpenStudy (anonymous):

but the minus roots doesnt count right? so its just, ( sqrt(4,5) , sqrt(2) )?

OpenStudy (anonymous):

i think they will all have the same value so just pick another point and see if it is bigger or smaller

OpenStudy (anonymous):

i need the second derivative and see if fxx is > 0 of < 0

OpenStudy (anonymous):

thank yoouu soo much coolsector <3

OpenStudy (anonymous):

youre welcome. i hope it is correct :)